# Superconducting pairing correlations on a trapped-ion quantum computer

### v2

@@ V Conclusion
First and foremost are the limits placed on adiabatic algorithms caused by entropy accumulation due to hardware errors. One option that appears sufficient (though maybe not necessary) to reach low-temperature states is to adiabatically evolve large systems for hundreds of hopping times [xu_neutral-atom_2025]. This is likely achievable with a factor 1000 more Trotter steps than we explored here [granet_dilution_2025, heyl_quantum_2019, mizuta_trotterization_2025, kovalsky_self-healing_2023], which translate to per-two-qubit-gate errors on the order of $10^{-6}$. While this requirement is mild compared to other quantum computing applications, understanding the most effective combination of (fermionic) error corrections codes [algaba_fermion--qubit_2025, chen_error-correcting_2024] and algorithmic techniques like bath-assisted cooling [matthies_programmable_2024, kishony_efficiently_2025] is essential to reach this goal. Reductions in gate error can also be used to simulate more complex Hamiltonians or longer evolution times which may be used to gain a better microscopic understanding of light-induced superconductivity in experimental setups [fausti_light-induced_2011, baykusheva_ultrafast_2022, fava_magnetic_2024].
On the technical side, we have introduced techniques to coherently inject states with locally fixed fermionic parity into fermionic encodings. These circuits can be used to access and explore superconducting phases and they show that ground state preparation for spin models can be useful beyond the study of magnetism. The techniques we have demonstrated can be adapted to other geometries, next-nearest neighbour hopping with a negative sign [xu_coexistence_2024], and other Hamiltonian terms, imparting, for example, the ability to introduce dopants coherently.
@@
One difficulty in measuring pairing correlations is that their local fermionic parity is odd when restricted to a single spin species. Previous Hubbard model simulations on quantum computers have used encodings in which the up- and down-fermions are encoded separately [arute_observation_2020, stanisic_observing_2022, hemery_measuring_2024, nigmatullin_experimental_2025]. On the one hand, this makes it relatively easy to find efficient circuits for the time evolution under the hopping terms of the Hamiltonian $t\sum_{\langle ij\rangle,\sigma}c^{\dagger}_{i\sigma}c_{j\sigma}$. On the other hand, in such encodings, measuring distant pairing correlations requires extended Jordan-Wigner strings, which makes it difficult to simultaneously measure many correlators, and leads to high-weight operators, which are highly susceptible to noise [granet_dilution_2025].
@@
To overcome these difficulties, in Figure S2, we show an “Octagon” fermion-to-qubit encoding. The encoding rules are closely related to the Compact Encoding [derby_compact_2021, derby2021compactfermionqubitmapping] and are as follows: Each site $i$ of the square-lattice Hubbard model is encoded by two qubits, one representing the “up” (blue, $\sigma=\uparrow$) and one representing the “down” (red, $\sigma=\downarrow$) fermionic modes. A diagonal edge is placed between each such pair. The square lattice is partitioned into even and odd plaquettes, corresponding to a checkerboard pattern in Fig S2a, with the plaquette at the south-west corner being an odd plaquette. Then, the odd plaquettes are further partitioned into $P$-plaquettes and $Q$-plaquettes, as illustrated in Fig S2a, so that plaquettes of one type contain disjoint sets of qubits. Edges are now placed between all nearest-neighbour “up”-qubits in the $P$-plaquettes and between all nearest-neighbour “down”-qubits in the $Q$-plaquettes. Depending on their geometry, these edges are horizontal or vertical edges. Within each odd plaquette, an ancilla qubits is placed. Thus each horizontal and vertical edge $(i\sigma,j\sigma)$ has a unique ancilla $a(i,j)$ associated to it. Edges are assigned an orientation, as shown via the arrows in Figure S2. Finally, identification between fermionic modes and qubit operators is made by associating to each edge $(i\sigma,j\sigma^{\prime})$ the Majorana bilinear $A_{i\sigma,j\sigma^{\prime}}=-i\gamma_{i\sigma}\gamma_{j\sigma^{\prime}}$, where $\gamma_{i\sigma}=c^{\dagger}_{i\sigma}+c_{i\sigma}$ and the qubit operator $A_{i\sigma,j\sigma^{\prime}}=X_{i\sigma}Y_{j\sigma^{\prime}}P$, where $P$ is $Y_{a(i,j)}$ for horizontal edges, $X_{a(i,j)}$ for down-pointing vertical edges, $X_{a(i,j)}$ for up-pointing vertical edges, and $\mathds{1}$ for diagonal edges. Note that the order of the indices of $A_{i\sigma,j\sigma^{\prime}}$ follows the orientation of arrows on the lattice (in this case $i\sigma\rightarrow j\sigma^{\prime})$, and reversed operators are defined as $A_{j\sigma^{\prime},i\sigma}=-A_{i\sigma,j\sigma^{\prime}}$. To complete the algebra of fermionic bilinears, we associate to each mode $i\sigma$ an on-site term $B_{i\sigma}=-i\gamma_{i\sigma}\overline{\gamma_{i\sigma}}\rightarrow Z_{i\sigma}$, where $\overline{\gamma_{i\sigma}}=i(c^{\dagger}_{i\sigma}-c_{i\sigma})$. Edge operators between non-adjacent modes $(i\sigma,k\sigma^{\prime})$ can be defined if there exists a path of edges $\{(j_{1},j_{2}),\dots(j_{l-1},j_{l})\}$, via $A_{i,k}=i^{l-1}\prod_{m=1}^{l-1}A_{j_{m},j_{m+1}}$, where $j_{1}=i\sigma$ and $j_{l}=k\sigma^{\prime}$. For paths $j_{1},\dots j_{l}$ that form closed loops $j_{l}=j_{1}$, one associates an operator $S=iA_{j_{1},j_{l}}$. The eigenvalues of $S=\pm 1$ indicate whether or not there is a $\pi$-magnetic flux through the enclosed area. If the flux-free sector is desired, one needs to ensure $S=1$ for all closed loops (this is the case for all topologically trivial loops of the experiments in this work, except for the closed loops that wrap around the torus, which are set to $\langle S\rangle=0$ to obtain mixed boundary conditions, see Section S13). Note that, in the qubit representation, the constraints on minimal loops that are topologically trivial are given by equations of the form $Z^{\otimes 8}XYXY=+1$, whereas if periodic boundary conditions were desired then one would need to set the qubit operators for minimal topologically non-trivial loops on the $6\times 6$ lattice (Figure S2) $Z^{\otimes 12}Y^{\otimes 6}=-1$ and $Z^{\otimes 12}X^{\otimes 6}=-1$, i.e., there is a negative sign for the non-trivial operators. The difference in sign arises due to the fact that the topologically trivial elementary loops touch an odd number of up-pointing vertical edges (namely, one), which are associated with a X-X operator on the ancilla.
@@
Figure S3: Octagon Fermion-to-Qubit Encoding for the operators and observables in the bilayer Hubbard Model. (a) Hamiltonian terms in the bilayer Hubbard model that is used to model bilayer nickelates [lu_interlayer-coupling-driven_2024, qu_bilayer_2024, schlomer_superconductivity_2024, oh_type-ii_2023]. The terms within each layer are the same as for in the single-layer case, namely a nearest-neighbour hopping $t\sum_{\langle ij\rangle\sigma}c^{\dagger}_{i\sigma}c_{j\sigma}$ and an on-site interaction $U\sum_{i}n_{i\uparrow}n_{i\downarrow}$. The layers are coupled via an exchange coupling $J\sum_{i}\mathbf{S}_{iA}\cdot\mathbf{S}_{iB}-n_{iA}n_{iB}/4$. (b) The system is mapped to two separate Octagon Encodings, one containing all “up”-modes of layer $B$ and all “down”-modes of layer AA (top) and one containing all “down”-modes of layer $B$ and all “up”-modes of layer AA (bottom). This is possible because all three Hamiltonian terms, as well as the rung-singlet-pairing observable $\Delta^{\dagger}_{AiBi}\Delta_{AjBj}$ have even fermionic parity in each subsystem.
@@
In Fig S6, we investigate the effect of noise on stabilisers and on local observables through numerical simulations. We simulate the circuit used for preparing the $(D_{\mathrm{Heisenberg}},D_{\mathrm{Hubbard}})=(1,1)$ state with a Pauli string decomposition, see Section S12.2. We study the effect of depolarizing noise on the circuit, at a fixed Pauli string truncation threshold $\epsilon=0.1/2^{15}$. We observe that both the $\eta$-pairing correlation and the average stabiliser expectation value undergo an exponential decay with the noise rate. This very smooth behaviour of expectation value with depolarizing noise is to be put in contrast with truncation error in classical numerical simulations techniques (see for example section S12). We see however that the stabilisers (which are weight $12$ Pauli strings) are damped significantly faster than the $\eta$-correlations (which are weight $4$). Such a strong effect of observable weight on noise sensitivity was observed and explained before [granet_dilution_2025]. We also compare these classical simulations with the values from the quantum computer for both $\eta$-pairing and the stabilisers. The hardware value is compatible with the classical depolarising noise simulations within two standard errors on the mean.
@@
The idea is shown in Fig. S7 and Fig. S8: The basic problem when Trotterising a sum of hopping terms on overlapping bonds is that an order must be chosen, i.e. one bond must be chosen before the other (alternatively particle number conservation can be broken [evered_probing_2025] which is unsuitable in this work since we rely on particle number conservation to measure pairing correlations). This breaking of time reversal-symmetry leads to error terms in the effective Hamiltonian that are currents from orbitals on the first bond to orbitals on the second bond e.g.,
@@
Note that this is a geometrically non-local circuit which allows us to prepare the toric code on the $3\times 3$ periodic ancilla lattice using a two-qubit depth of 4. This is the reason that we are not using a scheme based on measurement and feed-forward: While such schemes offer circuit depths that are constant in system size, for this particular system size, the minimal depth of 4 can be achieved unitarily with a smaller number of overall gates  [iqbal_topological_2024, fossfeig2023experimentaldemonstrationadvantageadaptive].
@@
this time with $\delta=-2/3$ exactly. The discussion of the bilayer perturbation theory is similar to the one in [schlomer_local_2024], although in that reference an exchange coupling between two layers of tJt-J-models was considered, whereas here we allow for the perturbation $V$ to create virtual states with doubly-occupied sites and only turn on the on-site interaction during the adiabatic evolution later on.
@@
The objective function to maximize is the overlap between the target state and the state generated by the circuit. The optimization over the circuit parameters is carried out through gradient descent over the infidelity. Specifically, we utilize the quasi-second order gradient descent method, L-BFGS algorithm, from SciPy [2020SciPy-NMeth]. With the 10-layer circuit, we obtain the fidelity $\mathcal{F}=0.99478$ with the best circuit realization from 30 random initializations. Notice each general $U(1)$ symmetry-preserving gate requires 3 elementary entangling 2-qubit gates, e.g., ZZPhase on our hardware. Due to a cancellation at the first layer, the total depth in terms of elemantary entangling 2-qubit gate of the XXZ state preparation circuit is 29. The parameters for all $10\times 4=40$ general $U(1)$-preserving gates used in the quantum experiments are given in Listing 7.
@@
To carry out the optimization of finding parameters of the fermionic circuits, here we utilize the Jordan-Wigner transformation for simplicity and efficiency. The Jordan-Wigner ordering we choose follows the qubit indices ordering. We first obtained both the $|s\rangle$ and $|d\rangle$ states with the Jordan-Wigner transformation. Then, we transform the variational fermionic circuit correspondingly according to the Jordan-Wigner ordering into a variational circuit over qubits by inserting fSWAPs. That is for any fermionic gate acting over not adjacent fermionic modes in the Jordan-Wigner ordering, we insert the fSWAPs. For example, the $U^{f}_{(1,3)}=\textrm{fSWAP}_{(1,2)}U_{(2,3)}\textrm{fSWAP}_{(1,2)}$. For a fermionic gate acting over adjacent fermionic modes, we apply no additional transformation. Now the problem reduces back to the standard variational quantum circuit optimization. We optimize the circuit with the quasi-second order gradient descent method, L-BFGS algorithm, from SciPy [2020SciPy-NMeth] over 30 different random initialization and select the best result. The best average fidelity is (2C)/2=0.99446(2-C)/2=0.99446.
@@
We use exact state vector simulation to obtain the exact energy gradient of all parameters. The optimization is carried out with the quasi-second order gradient descent method, L-BFGS algorithm, from SciPy [2020SciPy-NMeth] over 30 random initializations.
@@
We presented in the main text a benchmark of the hardware for the simulation of fermionic systems. The benchmark relied on the case $U=0$ where there is no interaction between the fermions, for which the imbalance can be computed in polynomial time. Such benchmark protocol was defined in Ref. [granet2025appqsim]. There, was introduced the notion of distinguishability cost to evaluate the quality of the output of a quantum hardware. This distinguishability cost is defined as the minimal amount of resources that an ideal perfect quantum computer has to spent to certify that the output of the tested hardware is incorrect. This amount of resources can be measured in shots or in number of gates to implement for example. In case of just one output $m$ with standard deviation $\sigma$ obtained from the hardware, and ideal value $m_{0}$ with standard deviation per shot $\sigma_{0}$ on the perfect hardware, this distinguishability cost is defined as
@@
where $N_{\rm gates}$ is the number of two-qubit gates in the circuit tested, and $\lceil\cdot\rceil$ the rounding to the nearest larger integer. This quantity $\delta$ corresponds exactly to the total number of two-qubit gates to implement (across different shots) on a perfect quantum computer to be confident by three standard deviations that the outcome of the tested hardware, whose precision is bounded by $\sigma$, is incorrect. The exact values $m_{0}$ and the variance per shot $\sigma_{0}$ can be computed exactly for this free-fermion benchmark. In case of multiple time points, the definition of the distinguishability cost is more involved and requires a $\chi^{2}$ test, see [granet2025appqsim]. The interpretation of this score is that before running $\delta$ two-qubit gates, the outcome of the tested hardware is statistically compatible with that of a perfect hardware. The score automatically balances shot noise and hardware noise and puts them on same footing.
@@
with $\sigma=1$. This filters the sum to only terms located at distance $d$ in Manhattan distance. We present the results in Fig S20c. We observe that as expected, the staggered $\eta$-pairing is strongly short-ranged right after the pulse, whereas after relaxation it vanishes at $d=0$, peaks at $d=2.3$ and has non-zero tails at larger distance. This increase of long-distance staggered $\eta$-pairing correlations after relaxation was also observed in a one-dimensional setting [kaneko_photoinduced_2019].
@@
The key question is then: how difficult, in terms of bond dimension is it to achieve order-one fidelities using the circuit-DMRG method, for the larger $6\times 6$ system? To answer this question, we show simulation fidelity as a function of bond dimension in Fig. S22b. Even the shallowest (2 steps) circuit achieves a fidelity roughly 3000 times lower that of the hardest Ising quench considered in [haghshenas_digital_2025] at bond dimension $\chi=2048$.
@@
The extremely small fidelity achieved for these circuits is not very surprising: A significant fraction of two-qubit gates are Clifford gates, most of which are used to couple the system qubits to the ancilla qubits during the fermionic hopping gadget. While these are necessary to enforce the fermion-to-qubit mapping, classical methods can achieve the correct fermionic statistics in other ways. For example, one can associate the physical degree of freedom of each MPS tensor to a fermionic mode and enforce the fermionic parity by introducing a $Z_{2}$ symmetry [cirac2021matrix]. To lower bound the cost of such an approach, we have also considered circuits in which all ancilla qubits (and gates acting on them) are removed, as well as all other gates that are responsible for the fermion-to-qubit encoding. The resulting simulation fidelities, shown in Fig. S22c, are much larger than those where the ancilla degrees of freedom are kept track of explicitly. In particular, the shallowest (2 steps) circuit now becomes comparable in difficulty to the $s=20$ Ising quench discussed above. As a result, we expect that a fermionic circuit-DMRG method utilizing additional symmetries, for example $SU(2)\times U(1)$ or $U(1)\times U(1)$ over spin and number of particles, may be able to capture the nearest-neighbour correlations accurately [liu2025accurate]. The situation on the deeper circuits (3 steps and 4 steps), where after the light pulse additional Hubbard-Trotter steps are applied to relax the state, remain out of reach.
@@
In addition to the circuit-DMRG results obtained from the FermioniQ Ava emulator, we have also run some simulations of the circuit written with the Jordan-Wigner (JW) encoding, alleviating the difficulty of simulating the compact encoding, at the price of adding long range JW strings. The JW string order is chosen to match the ordering of MPS tensors, making this approach similar to the fermionic MPS approach without having to utilize an extra $Z_{2}$ symmetry. These MPS simulations were performed using a slightly simpler approach where quantum gates are contracted into the MPS one at a time. The algorithm supports native application of two-qubit gates between non-adjacent qubits, as well as multi-qubit gates of the form $\exp(i\theta P)$ for arbitrary Pauli strings. These are applied by first obtaining their exact representation as an MPO of bond dimension 2, which is then contracted into the MPS following the zip-up algorithm from [Stoudenmire_2010], with one caveat: the zip-up algorithm is carried out using QR decompositions, rather than SVD. After the gate has been contracted into the MPS, the latter is put in canonical form with the center at the leftmost bond affected by the gate, followed by SVD truncation down to the target bond dimension; this is repeated on all affected bonds from left to right. As discussed in [Zhou2020], the sum of the squared singular values we keep determines the fidelity of each single truncation, and the product of all of these is a standard estimate of the fidelity of the final state. The algorithm runs on GPUs, implemented using NVIDIA cuTensorNet and is publicly available at [PabloAndresCQ] under the name MPSxGate. The results can be seen (blue crosses) in Fig. S22c. Interestingly, the fidelity match closely the circuit simulations where all ancilla have been removed. This indicates the naive estimate by removing gates related to ancilla gives a reasonable estimate of the complexity of simulating with fermionic MPS.

### v3

@@ V Conclusion
First and foremost are the limits placed on adiabatic algorithms caused by entropy accumulation due to hardware errors. One option that appears sufficient (though maybe not necessary) to reach low-temperature states is to adiabatically evolve large systems for hundreds of hopping times [84]. This is likely achievable with a factor 1000 more Trotter steps than we explored here [35, 39, 58, 44], which translate to per-two-qubit-gate errors on the order of $10^{-6}$. While this requirement is mild compared to other quantum computing applications, understanding the most effective combination of (fermionic) error corrections codes [1, 20] and algorithmic techniques like bath-assisted cooling [55, 43] is essential to reach this goal. Reductions in gate error can also be used to simulate more complex Hamiltonians or longer evolution times which may be used to gain a better microscopic understanding of light-induced superconductivity in experimental setups [30, 10, 31].
On the technical side, we have introduced techniques to coherently inject states with locally fixed fermionic parity into fermionic encodings. These circuits can be used to access and explore superconducting phases and they show that ground state preparation for spin models can be useful beyond the study of magnetism. The techniques we have demonstrated can be adapted to other geometries, next-nearest neighbour hopping with a negative sign [83], and other Hamiltonian terms, imparting, for example, the ability to introduce dopants coherently.
@@
One difficulty in measuring pairing correlations is that their local fermionic parity is odd when restricted to a single spin species. Previous Hubbard model simulations on quantum computers have used encodings in which the up- and down-fermions are encoded separately [7, 73, 38, 59]. On the one hand, this makes it relatively easy to find efficient circuits for the time evolution under the hopping terms of the Hamiltonian $t\sum_{\langle ij\rangle,\sigma}c^{\dagger}_{i\sigma}c_{j\sigma}$. On the other hand, in such encodings, measuring distant pairing correlations requires extended Jordan-Wigner strings, which makes it difficult to simultaneously measure many correlators, and leads to high-weight operators, which are highly susceptible to noise [35].
@@
To overcome these difficulties, in Figure S2, we show an “Octagon” fermion-to-qubit encoding. The encoding rules are closely related to the Compact Encoding [24, 25] and are as follows: Each site $i$ of the square-lattice Hubbard model is encoded by two qubits, one representing the “up” (blue, $\sigma=\uparrow$) and one representing the “down” (red, $\sigma=\downarrow$) fermionic modes. A diagonal edge is placed between each such pair. The square lattice is partitioned into even and odd plaquettes, corresponding to a checkerboard pattern in Fig S2a, with the plaquette at the south-west corner being an odd plaquette. Then, the odd plaquettes are further partitioned into $P$-plaquettes and $Q$-plaquettes, as illustrated in Fig S2a, so that plaquettes of one type contain disjoint sets of qubits. Edges are now placed between all nearest-neighbour “up”-qubits in the $P$-plaquettes and between all nearest-neighbour “down”-qubits in the $Q$-plaquettes. Depending on their geometry, these edges are horizontal or vertical edges. Within each odd plaquette, an ancilla qubits is placed. Thus each horizontal and vertical edge $(i\sigma,j\sigma)$ has a unique ancilla $a(i,j)$ associated to it. Edges are assigned an orientation, as shown via the arrows in Figure S2. Finally, identification between fermionic modes and qubit operators is made by associating to each edge $(i\sigma,j\sigma^{\prime})$ the Majorana bilinear $A_{i\sigma,j\sigma^{\prime}}=-i\gamma_{i\sigma}\gamma_{j\sigma^{\prime}}$, where $\gamma_{i\sigma}=c^{\dagger}_{i\sigma}+c_{i\sigma}$ and the qubit operator $A_{i\sigma,j\sigma^{\prime}}=X_{i\sigma}Y_{j\sigma^{\prime}}P$, where $P$ is $Y_{a(i,j)}$ for horizontal edges, $X_{a(i,j)}$ for down-pointing vertical edges, $X_{a(i,j)}$ for up-pointing vertical edges, and $\mathds{1}$ for diagonal edges. Note that the order of the indices of $A_{i\sigma,j\sigma^{\prime}}$ follows the orientation of arrows on the lattice (in this case $i\sigma\rightarrow j\sigma^{\prime})$, and reversed operators are defined as $A_{j\sigma^{\prime},i\sigma}=-A_{i\sigma,j\sigma^{\prime}}$. To complete the algebra of fermionic bilinears, we associate to each mode $i\sigma$ an on-site term $B_{i\sigma}=-i\gamma_{i\sigma}\overline{\gamma_{i\sigma}}\rightarrow Z_{i\sigma}$, where $\overline{\gamma_{i\sigma}}=i(c^{\dagger}_{i\sigma}-c_{i\sigma})$. Edge operators between non-adjacent modes $(i\sigma,k\sigma^{\prime})$ can be defined if there exists a path of edges $\{(j_{1},j_{2}),\dots(j_{l-1},j_{l})\}$, via $A_{i,k}=i^{l-1}\prod_{m=1}^{l-1}A_{j_{m},j_{m+1}}$, where $j_{1}=i\sigma$ and $j_{l}=k\sigma^{\prime}$. For paths $j_{1},\dots j_{l}$ that form closed loops $j_{l}=j_{1}$, one associates an operator $S=iA_{j_{1},j_{l}}$. The eigenvalues of $S=\pm 1$ indicate whether or not there is a $\pi$-magnetic flux through the enclosed area. If the flux-free sector is desired, one needs to ensure $S=1$ for all closed loops (this is the case for all topologically trivial loops of the experiments in this work, except for the closed loops that wrap around the torus, which are set to $\langle S\rangle=0$ to obtain mixed boundary conditions, see Section S13). Note that, in the qubit representation, the constraints on minimal loops that are topologically trivial are given by equations of the form $Z^{\otimes 8}XYXY=+1$, whereas if periodic boundary conditions were desired then one would need to set the qubit operators for minimal topologically non-trivial loops on the $6\times 6$ lattice (Figure S2) $Z^{\otimes 12}Y^{\otimes 6}=-1$ and $Z^{\otimes 12}X^{\otimes 6}=-1$, i.e., there is a negative sign for the non-trivial operators. The difference in sign arises due to the fact that the topologically trivial elementary loops touch an odd number of up-pointing vertical edges (namely, one), which are associated with a X-X operator on the ancilla.
@@
Figure S3: Octagon Fermion-to-Qubit Encoding for the operators and observables in the bilayer Hubbard Model. (a) Hamiltonian terms in the bilayer Hubbard model that is used to model bilayer nickelates [52, 65, 72, 60]. The terms within each layer are the same as for in the single-layer case, namely a nearest-neighbour hopping $t\sum_{\langle ij\rangle\sigma}c^{\dagger}_{i\sigma}c_{j\sigma}$ and an on-site interaction $U\sum_{i}n_{i\uparrow}n_{i\downarrow}$. The layers are coupled via an exchange coupling $J\sum_{i}\mathbf{S}_{iA}\cdot\mathbf{S}_{iB}-n_{iA}n_{iB}/4$. (b) The system is mapped to two separate Octagon Encodings, one containing all “up”-modes of layer $B$ and all “down”-modes of layer AA (top) and one containing all “down”-modes of layer $B$ and all “up”-modes of layer AA (bottom). This is possible because all three Hamiltonian terms, as well as the rung-singlet-pairing observable $\Delta^{\dagger}_{AiBi}\Delta_{AjBj}$ have even fermionic parity in each subsystem.
@@
In Fig S6, we investigate the effect of noise on stabilisers and on local observables through numerical simulations. We simulate the circuit used for preparing the $(D_{\mathrm{Heisenberg}},D_{\mathrm{Hubbard}})=(1,1)$ state with a Pauli string decomposition, see Section S12.2. We study the effect of depolarizing noise on the circuit, at a fixed Pauli string truncation threshold $\epsilon=0.1/2^{15}$. We observe that both the $\eta$-pairing correlation and the average stabiliser expectation value undergo an exponential decay with the noise rate. This very smooth behaviour of expectation value with depolarizing noise is to be put in contrast with truncation error in classical numerical simulations techniques (see for example section S12). We see however that the stabilisers (which are weight $12$ Pauli strings) are damped significantly faster than the $\eta$-correlations (which are weight $4$). Such a strong effect of observable weight on noise sensitivity was observed and explained before [35]. We also compare these classical simulations with the values from the quantum computer for both $\eta$-pairing and the stabilisers. The hardware value is compatible with the classical depolarising noise simulations within two standard errors on the mean.
@@
The idea is shown in Fig. S7 and Fig. S8: The basic problem when Trotterising a sum of hopping terms on overlapping bonds is that an order must be chosen, i.e. one bond must be chosen before the other (alternatively particle number conservation can be broken [29] which is unsuitable in this work since we rely on particle number conservation to measure pairing correlations). This breaking of time reversal-symmetry leads to error terms in the effective Hamiltonian that are currents from orbitals on the first bond to orbitals on the second bond e.g.,
@@
Note that this is a geometrically non-local circuit which allows us to prepare the toric code on the $3\times 3$ periodic ancilla lattice using a two-qubit depth of 4. This is the reason that we are not using a scheme based on measurement and feed-forward: While such schemes offer circuit depths that are constant in system size, for this particular system size, the minimal depth of 4 can be achieved unitarily with a smaller number of overall gates  [40, 32].
@@
this time with $\delta=-2/3$ exactly. The discussion of the bilayer perturbation theory is similar to the one in [71], although in that reference an exchange coupling between two layers of tJt-J-models was considered, whereas here we allow for the perturbation $V$ to create virtual states with doubly-occupied sites and only turn on the on-site interaction during the adiabatic evolution later on.
@@
The objective function to maximize is the overlap between the target state and the state generated by the circuit. The optimization over the circuit parameters is carried out through gradient descent over the infidelity. Specifically, we utilize the quasi-second order gradient descent method, L-BFGS algorithm, from SciPy [80]. With the 10-layer circuit, we obtain the fidelity $\mathcal{F}=0.99478$ with the best circuit realization from 30 random initializations. Notice each general $U(1)$ symmetry-preserving gate requires 3 elementary entangling 2-qubit gates, e.g., ZZPhase on our hardware. Due to a cancellation at the first layer, the total depth in terms of elemantary entangling 2-qubit gate of the XXZ state preparation circuit is 29. The parameters for all $10\times 4=40$ general $U(1)$-preserving gates used in the quantum experiments are given in Listing 7.
@@
To carry out the optimization of finding parameters of the fermionic circuits, here we utilize the Jordan-Wigner transformation for simplicity and efficiency. The Jordan-Wigner ordering we choose follows the qubit indices ordering. We first obtained both the $|s\rangle$ and $|d\rangle$ states with the Jordan-Wigner transformation. Then, we transform the variational fermionic circuit correspondingly according to the Jordan-Wigner ordering into a variational circuit over qubits by inserting fSWAPs. That is for any fermionic gate acting over not adjacent fermionic modes in the Jordan-Wigner ordering, we insert the fSWAPs. For example, the $U^{f}_{(1,3)}=\textrm{fSWAP}_{(1,2)}U_{(2,3)}\textrm{fSWAP}_{(1,2)}$. For a fermionic gate acting over adjacent fermionic modes, we apply no additional transformation. Now the problem reduces back to the standard variational quantum circuit optimization. We optimize the circuit with the quasi-second order gradient descent method, L-BFGS algorithm, from SciPy [80] over 30 different random initialization and select the best result. The best average fidelity is (2C)/2=0.99446(2-C)/2=0.99446.
@@
We use exact state vector simulation to obtain the exact energy gradient of all parameters. The optimization is carried out with the quasi-second order gradient descent method, L-BFGS algorithm, from SciPy [80] over 30 random initializations.
@@
We presented in the main text a benchmark of the hardware for the simulation of fermionic systems. The benchmark relied on the case $U=0$ where there is no interaction between the fermions, for which the imbalance can be computed in polynomial time. Such benchmark protocol was defined in Ref. [34]. There, was introduced the notion of distinguishability cost to evaluate the quality of the output of a quantum hardware. This distinguishability cost is defined as the minimal amount of resources that an ideal perfect quantum computer has to spent to certify that the output of the tested hardware is incorrect. This amount of resources can be measured in shots or in number of gates to implement for example. In case of just one output $m$ with standard deviation $\sigma$ obtained from the hardware, and ideal value $m_{0}$ with standard deviation per shot $\sigma_{0}$ on the perfect hardware, this distinguishability cost is defined as
@@
where $N_{\rm gates}$ is the number of two-qubit gates in the circuit tested, and $\lceil\cdot\rceil$ the rounding to the nearest larger integer. This quantity $\delta$ corresponds exactly to the total number of two-qubit gates to implement (across different shots) on a perfect quantum computer to be confident by three standard deviations that the outcome of the tested hardware, whose precision is bounded by $\sigma$, is incorrect. The exact values $m_{0}$ and the variance per shot $\sigma_{0}$ can be computed exactly for this free-fermion benchmark. In case of multiple time points, the definition of the distinguishability cost is more involved and requires a $\chi^{2}$ test, see [34]. The interpretation of this score is that before running $\delta$ two-qubit gates, the outcome of the tested hardware is statistically compatible with that of a perfect hardware. The score automatically balances shot noise and hardware noise and puts them on same footing.
@@
with $\sigma=1$. This filters the sum to only terms located at distance $d$ in Manhattan distance. We present the results in Fig S20c. We observe that as expected, the staggered $\eta$-pairing is strongly short-ranged right after the pulse, whereas after relaxation it vanishes at $d=0$, peaks at $d=2.3$ and has non-zero tails at larger distance. This increase of long-distance staggered $\eta$-pairing correlations after relaxation was also observed in a one-dimensional setting [41].
@@
The key question is then: how difficult, in terms of bond dimension is it to achieve order-one fidelities using the circuit-DMRG method, for the larger $6\times 6$ system? To answer this question, we show simulation fidelity as a function of bond dimension in Fig. S22b. Even the shallowest (2 steps) circuit achieves a fidelity roughly 3000 times lower that of the hardest Ising quench considered in [37] at bond dimension $\chi=2048$.
@@
The extremely small fidelity achieved for these circuits is not very surprising: A significant fraction of two-qubit gates are Clifford gates, most of which are used to couple the system qubits to the ancilla qubits during the fermionic hopping gadget. While these are necessary to enforce the fermion-to-qubit mapping, classical methods can achieve the correct fermionic statistics in other ways. For example, one can associate the physical degree of freedom of each MPS tensor to a fermionic mode and enforce the fermionic parity by introducing a $Z_{2}$ symmetry [21]. To lower bound the cost of such an approach, we have also considered circuits in which all ancilla qubits (and gates acting on them) are removed, as well as all other gates that are responsible for the fermion-to-qubit encoding. The resulting simulation fidelities, shown in Fig. S22c, are much larger than those where the ancilla degrees of freedom are kept track of explicitly. In particular, the shallowest (2 steps) circuit now becomes comparable in difficulty to the $s=20$ Ising quench discussed above. As a result, we expect that a fermionic circuit-DMRG method utilizing additional symmetries, for example $SU(2)\times U(1)$ or $U(1)\times U(1)$ over spin and number of particles, may be able to capture the nearest-neighbour correlations accurately [51]. The situation on the deeper circuits (3 steps and 4 steps), where after the light pulse additional Hubbard-Trotter steps are applied to relax the state, remain out of reach.
@@
In addition to the circuit-DMRG results obtained from the FermioniQ Ava emulator, we have also run some simulations of the circuit written with the Jordan-Wigner (JW) encoding, alleviating the difficulty of simulating the compact encoding, at the price of adding long range JW strings. The JW string order is chosen to match the ordering of MPS tensors, making this approach similar to the fermionic MPS approach without having to utilize an extra $Z_{2}$ symmetry. These MPS simulations were performed using a slightly simpler approach where quantum gates are contracted into the MPS one at a time. The algorithm supports native application of two-qubit gates between non-adjacent qubits, as well as multi-qubit gates of the form $\exp(i\theta P)$ for arbitrary Pauli strings. These are applied by first obtaining their exact representation as an MPO of bond dimension 2, which is then contracted into the MPS following the zip-up algorithm from [74], with one caveat: the zip-up algorithm is carried out using QR decompositions, rather than SVD. After the gate has been contracted into the MPS, the latter is put in canonical form with the center at the leftmost bond affected by the gate, followed by SVD truncation down to the target bond dimension; this is repeated on all affected bonds from left to right. As discussed in [88], the sum of the squared singular values we keep determines the fidelity of each single truncation, and the product of all of these is a standard estimate of the fidelity of the final state. The algorithm runs on GPUs, implemented using NVIDIA cuTensorNet and is publicly available at [4] under the name MPSxGate. The results can be seen (blue crosses) in Fig. S22c. Interestingly, the fidelity match closely the circuit simulations where all ancilla have been removed. This indicates the naive estimate by removing gates related to ancilla gives a reasonable estimate of the complexity of simulating with fermionic MPS.